OpenStudy (azionne):

4(7 - 3x) = 7(4 - 2x) will medal

pooja195 (pooja195):

Start by distributing \[\rm~(4)(7)+(4)(−3x)=(7)(4)+(7)(−2x)\]

pooja195 (pooja195):



OpenStudy (igigighjkl):

do you want awnser?

OpenStudy (igigighjkl):

O.k I won't.

OpenStudy (pawanyadav):

Apply everything that you got from all previous question:-

OpenStudy (azionne):

i do but pooja wont approve and nether will i.

pooja195 (pooja195):

=) Thats more like it! Alright right the equation that you get once you distribute :)

OpenStudy (azionne):

28-12x=28-14x

pooja195 (pooja195):

−12x+28=−14x+28 Good Add 14x to both sides. \[~\rm~−12x+28+14x=−14x+28+14x\]

OpenStudy (igigighjkl):

MAN THERE IS NOTHING I CAN EVER AWNSER OR HELP WITH!!!!!! ;(

OpenStudy (azionne):

2+28=28

OpenStudy (azionne):

is that right?

OpenStudy (pawanyadav):

Wow you learnt- you can also do it to the answer ..otherwise we are here

pooja195 (pooja195):

2x+28=28 Subtract 28 from both sides. \[\huge~\rm~2x+28−28=28−28\]

OpenStudy (azionne):

i got x=0

pooja195 (pooja195):

Well done correct! :)

OpenStudy (anonymous):

im glad for you @Azionne :) keep learning

OpenStudy (azionne):

^ i will

pooja195 (pooja195):

=)

OpenStudy (anonymous):

All the best :)

OpenStudy (anonymous):

\(\Huge\ddot\smile\)

OpenStudy (pawanyadav):

I know you can do it

OpenStudy (azionne):

lol yes i can

pooja195 (pooja195):

That's the spirit :)

OpenStudy (chantysquirrel1129**):

dang, I wish I had that self esteem xD

OpenStudy (chantysquirrel1129**):

But good job dewd :D